OpenStudy (anonymous):

Solve. x=sqr rt x+13+7

OpenStudy (anonymous):

is it \[x=\sqrt{x+13}+7\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if so then \[x-7=\sqrt{x+13}\] square both sides

OpenStudy (anonymous):

\[x^2-14x+49=x+13\] \[x^2-15x+36=0\]

OpenStudy (anonymous):

factor \[(x-13)(x-2)=0\] \[x=13\] \[x=2\]

OpenStudy (anonymous):

now you have to check because step two was squaring.

OpenStudy (anonymous):

oops i factored wrong!

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

\[(x-12)(x-3)=0\] \[x=12\] \[x=3\]

OpenStudy (anonymous):

now check and you will see that 12 works and 3 does not

OpenStudy (anonymous):

ok. thats what i got. thanks for confirming. medal

OpenStudy (anonymous):

so answer is 12

OpenStudy (anonymous):

welcome